A sewing machine needle moves up and down in simple harmonic motion with an amplitude of 1.27cm and a frequency of 2.55Hz. what is the maximum speed of the needle.
Answer #1

Solution) Amplitude , A = 1.27 cm = 1.27×10^(-2) m

Frequency , f = 2.55 Hz

Maximum speed , V = ?

V = Aw

But w = 2(pi)f = 2(pi)×2.55 = 16.02 rad/s

V = Aw

V = 1.27×10^(-2)×16.02

V = 0.203 m/s